From 160 gm of SO2 sample 1.20×10^24 molecules are removed then the volume of left over SO2 AT S.T.P WILL BE???

Grade:11

1 Answers

achhita ranjan
5 years ago
Mole of 160 gm of so2 is 2.5
No of molecules in 2.5 mole of so2=1.505×10^24
No of molecules left -1.505×10^24-1.20×10^24=
                                        =1.305×10^24
Mole of left so2=1.305 ×10^24/6.022×10^24
                               =2.167 approx
Volume=2.167×22.4=48.54L